OpenStudy (anonymous):

A card is selected from a standard deck of 52 playing cards. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). Find the probability of selecting · a five given the card is a not a club. · a heart given the card is red. · a face card, given that the card is black. Show step by step work. Give all solutions exactly in reduced fraction form.

OpenStudy (anonymous):

the condition "is not a club" is unimportant. you have 12 face cards and four aces???

OpenStudy (anonymous):

if it really has 12 face cards then there is no 5. what kind of trick deck is this?

OpenStudy (anonymous):

if it is really a standard deck then there are 13 cards of each denomination, one is a 5 and the probability you pick a 5 is \[\frac{1}{13}\]

OpenStudy (anonymous):

of the red cards, half are diamonds and half are hearts so probability you pick a heart is \[\frac{1}{2}\]

OpenStudy (anonymous):

out of each suit of 13 cards 3 are face cards so probability you get a face card is \[\frac{3}{13}\]

OpenStudy (anonymous):

probability of selecting a five given the card is a not a club clubs,hearts,spades,diamonds there are 4 fives in standard deck so, 3/52

OpenStudy (anonymous):

HOWEVER this line A standard deck of cards has 12 face cards and four Aces makes no sense at all

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

wouldnt it just be 3/(54-4)

OpenStudy (anonymous):

so 3/50

OpenStudy (anonymous):

@hahd why 54-4

OpenStudy (anonymous):

oh crap sorry didnt read question properly

OpenStudy (anonymous):

i mean 52-13

OpenStudy (anonymous):

sorry just wokeup

OpenStudy (anonymous):

cause all of clubs are gone

OpenStudy (anonymous):

1.probability of selecting a five given the card is a not a club clubs,hearts,spades,diamonds there are 4 fives in standard deck so, 3/52 2. a heart given the card is red. there are 13 hearts in a deck so, 13/52=1/4 3.a face card, given that the card is black. there are 12 face cards in a deck and 6 of them are black so, 6/52=3/26

OpenStudy (anonymous):

what is 3/52 = to

OpenStudy (anonymous):

its 3/52